OpenStudy (anonymous):

What is h^-1(-4x+7) if h(x) = 3 -2x? My answer is 2(x-1), is it correct?

OpenStudy (anonymous):

let y=3-2x. 2x=3-y x=(3-y)/2 So, inverse of f is :(3-x)/2 substitute -4x+7 i this.we get (3+4x-7)/2 =2(x-1). yeah,your answer is right.